# The HITRAN2024 methane update

### v1

@@ Abstract
Spectroscopic parameters of methane from many different studies were gathered to improve the HITRAN database towards its 2024 version. After a validation process using high-resolution FTS and CRDS spectra, about 60,000 lines of the four most abundant isotopologues were replaced from the dyad to the triacontad regions. These changes amount to $56,000$ transition wavenumbers, $16,000$ line intensities, $30,000$ pressure-broadening half-widths, and 3500 assignments. 40,000 new lines were added, extending the database from 12,000 cm-1 up to 14,000 cm-1, and covering some gaps. A greater focus was brought on the pentad, octad, and tetradecad regions, targeted by several remote instruments. In these regions, comparisons of spectral fits from multiple line lists were performed, taking only the parameters that provide best fit for each line. In the $\nu_{3}$ band, in addition to replacing the previous values, gathered speed-independent pressure broadenings of 12CH4 were used to fit Padé-approximant functions. These functions then replaced any outdated experimental data in $\nu_{3}$, missing data in the new lines, as well as the values that were determined to be outside their physical boundaries. The CH3D broadenings were replaced in the same manner, for missing and low or high values, using a semi-empirical formula instead.
@@ 1 Introduction
The octad (3500–5000 cm-1) and tetradecad regions (5000–6250 cm-1) in particular are crucial for satellite missions [lorente_methane_2021, malina_consistency_2022, chesnokova_estimation_2020], not only for its principal isotopologue but also for 13CH4 and CH3D, which are used to determine 13C/12C and D/H ratios. The octad has been improved in previous editions of HITRAN in the window observed by TROPOMI (4190–4340 cm-1). However, at higher wavenumbers, spectral parameters were of lower accuracy, incomplete, or both. Current and planned CH4 satellite sensors strive for  0.5% precision in the column (Ref. [jacob_quantifying_2022]; Table 1) for area flux mapping purposes, which requires similar accuracies for the intensities and widths and is not achieved in many regions. Even in the TROPOMI window, the HITRAN uncertainty code only reaches 4 (i.e., $0.0001\leq\Delta\nu<0.001$ cm-1 for line positions and $10\%\leq\frac{\Delta S_{ij}}{S_{ij}}<20\%$ for intensities). Above, and including the tetradecad region, other problems arise from the lack of assignment and/or value for the lower state energy ($E^{\prime\prime}$). Given the temperature range, these issues are limited in the Earth’s atmosphere, but make accurate spectral modeling impossible at more extreme temperatures (for example, low temperatures on Titan or high temperatures on hot-Jupiters and brown dwarfs). Furthermore, the origin of some of the parameters present in HITRAN2020 is not well documented, especially for the pressure-induced data. This can cause additional problems, preventing comparisons or adjustments in the case of models.
@@
The focus of the HITRAN2024 update was on transition wavenumbers, intensities, line-shape parameters due to pressure of air, and assignments. Significant progress has been made between 900 and 14000 cm-1 [kefala_empirical_2024, nikitin_improved_2024, rodina_improved_2021, nikitin_preliminary_2013, starikova_assignment_2024, malarich_dual_2021, nikitin_first_2024, campargue_high_2023] for the three most abundant isotopologues, 12CH4, 13CH4, and 12CH3D, allowing for an update to the transition wavenumber, intensity, and assignments from the HITRAN2020 database using various recorded spectra for validation. These line lists also cover missing portions of the infrared spectrum that were previously challenging to obtain, as well as some in-between lines that were missing. Similar efforts have been put into obtaining good quality pressure-induced line-shape parameters [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023, dudaryonok_semiempirical_2018], with the differences being that we considered older work, with the intention of replacing the parameters for more lines than those in Section 2 to improve traceability. To this end, a combination of good quality measurements, with an empirical Padé-approximants model based on these measurements and semi-empirical models to fill the gaps was used. These results provide the the necessary foundation for the inclusion of first-order and full line-mixing with speed-dependent profiles.
2 Additions and updates
@@
The newest versions of MARVELized-ExoMol [yurchenko_exomol_2024], TheoReTS [10.3847/1538-4357/aa8909], HITEMP [10.3847/1538-4365/ab7a1a] and MeCaSDa [MeCaSDa_2013], which is partly accessible via VAMDC portal [rixon_VAMDC_2011, albert_VAMDC_2020], line lists as well as lists from Nikitin et al. [nikitin_improved_2024, nikitin_preliminary_2013, nikitin_first_2024], Rodina et al. [rodina_improved_2021], Sung et al. [sung_new_2024], Starikova et al. [starikova_assignment_2024], Malarich et al. [malarich_dual_2021], and Campargue et al. [campargue_high_2023] were available for the HITRAN2024 update and are summarized in Sec. 2.1. These works were then validated against several high-quality, high-resolution FTS [2015ApJ...813...12H, bertin_co2_2024, birk_esa_2017, birk_measurement_2025] and CRDS spectra [birk_esa_2017] obtained under well-controlled laboratory conditions. The best parameters depending on the spectral region were selected and added to the HITRAN2024 database.
@@
Much of the focus was brought to the pentad, octad, and tetradecad regions, but many other spectral regions have also been updated for HITRAN2024. Changes from these lists were applied almost exclusively for the transition wavenumbers, intensities, and energy levels, pressure-induced parameters of most of the available line lists were missing or approximate. Updates on these parameters are detailed in Sec. 3. The newly added lines, which were formerly missing in HITRAN, were added as is, but their pressure broadenings and shifts were replaced as well.
2.1 New line lists at disposal
This section summarizes recent studies for methane infrared spectroscopy. Each line list was subjected to a validation process to determine the parameters to include in the HITRAN2024 update. The specific additions from these works are described in Sec. 2.2.
@@ 2.1.1 ExoMol
The ExoMol project [tennyson_2024_2024] has produced a number of methane line lists of increasing completeness and accuracy designed for use at high temperatures [yurchenko_exomol_2014, yurchenko_hybrid_2017, yurchenko_exomol_2024]. The intensities in this list are calculated variationally. The line positions in the most recent of these [yurchenko_exomol_2024], known as the MM line list, was based on a comprehensive MARVEL (measured active vibration-rotation energy levels) study of 12CH4 by Kefala et al [kefala_empirical_2024]. This MARVEL study extracted 82173 rovibrational and rotational transitions from 96 literature sources containing high-resolution studies of 12CH4 spectra; the study identified a further 42 sources containing unassigned or partially assigned methane spectra which could not be included in the MARVEL study. An analysis of these sources is currently in progress and has so far led to the assignment of almost 10000 further lines, which extend the coverage of the previous study in terms of coverage of both rotational and vibrational states. This work can enable updates to future releases of both HITRAN and HITEMP. A HITRAN-format 296 K version of the MM line list generated using pyExoCross [Zhang_pyExoCross_2024] was used in this study.
@@ 2.1.5 IAO FTIR measurements 900–1050 cm-1
2.1.6 Kitt Peak & JPL FTS measurements of 13CH4 4000–4700 cm-1
Multiple sets of 13CH4 line lists, either empirical or model predictions, were available from various sources, e.g., HITRAN 2020 [gordon_hitran2020_2022, brown_measurements_2016], JPL GFIT package (“atm.161”) [toon_spectrometric_2021], and Kitt Peak FTS (291 K) and JPl-FTS (80 K). More importantly, comprehensive updates to the previous work by Brown et al. [brown_measurements_2016] became available through the effective Hamiltonian modeling extended to tetradecad region for hot band transitions. All of the line lists have been evaluated and cross-compared against high-resolution laboratory spectra obtained at 291 K and 80 K, by which their quantum identifications and lower state energy values have been determined as presented in Sung et al. [sung_new_2024]
@@ 2.1.6 Kitt Peak & JPL FTS measurements of 13CH4 4000–4700 cm-1
A preliminary version of the composite line list has been implemented, followed by a final version to be available on-line, in which more quantum assignments will be available. Table 1 presents statistics on the improvements (both additions and updates inclusive) as reported by Sung et al. [sung_new_2024], showing an almost ten-fold increase in the total number of line list entries in the 4000–4700 cm-1 wavenumber interval for 13CH4, when compared to the HITRAN2020.
@@
Figure 1: Comparison of synthetic spectra generated with HITRAN2020 (in pink) and the updated composite line list, our update (in red) with high-resolution Kitt Peak FTS spectrum at 291 K (Left, in black) and JPL-FTS spectrum at 80 K (Right, in blue), showing better performance and spectral coverage by the new composite line list. Note that the synthetic spectra were computed under the same conditions as the corresponding laboratory spectra. Any H2O (not presented) and 12CH4 features (in green) were identified during the line-by-line matching, as presented for 12CH4 features.
@@
Finally, Fig. 1 presents a graphical summary of the updates, showing synthetic spectra generated with the HITRAN2020 line list and the composite line list reported in this work compared to two high-resolution laboratory FTIR spectra, respectively at 291 K( Left panel) and 80 K (Right panel). The synthetic spectra at 291 K and 80 K, respectively, were computed under the same conditions as the corresponding laboratory spectra. It is shown that the composite line list covers much wider spectral features across the Octad region as well as provides a better representation of the observed spectra at both temperatures thanks to the lower-state energy updates.
2.1.7 IAO analyses of FTS measurements 4970–6200 cm-1
@@
2.1.8 NIST, DLR & USTC measurements 5970–6250 cm-1
The tetradecad spectral region has been the subject of extensive study, including the WKLMC empirical lineline [campargue_2013], the GOSAT 2014 line list [nikitin_gosat-2014_2015] (which was the basis of both HITRAN2016 and HITRAN2020 for this spectral region), the widely used empirical line list of Devi 2015 [devi_self-_2015], and the Nikitin 2017 analysis of WKLMC [nikitin_analysis_2017]. Significant differences exist between these line lists, complicated by correlation between parameters in these dense and complex spectra and the difficulty in identifying and assigning weaker transitions. These complications lead to line lists that trade parameter accuracy for self-consistency, thus impeding efforts to refine line parameters.

### v2

@@ Abstract
Spectroscopic parameters of methane from many different studies were gathered to improve the HITRAN database towards its 2024 version. After a validation process using high-resolution FTS and CRDS spectra, about 80,000 lines of the four most abundant isotopologues were replaced from the dyad to the triacontad regions. These changes amount to 51,000 transition wavenumbers, 18,000 line intensities, 33,000 pressure-broadening half-widths, and 3300 assignments. 44,000 new lines were added with 16,000 old lines removed, extending the database from 12,000 cm-1 up to 14,000 cm-1, and covering some gaps. A greater focus was brought on the pentad, octad, and tetradecad regions, targeted by several remote sensing instruments. In these regions, comparisons of spectral fits from multiple line lists were performed, taking only the parameters that provide best fit for each line. In the $\nu_{3}$ band, in addition to replacing the previous values, speed-independent pressure broadening parameters of 12CH4 were gathered and used to fit Padé-approximant functions. These functions then replaced any outdated experimental data in $\nu_{3}$, missing data in the new lines, as well as the values that were determined to be outside their physical boundaries. The CH3D broadening parameters were replaced in the same manner, for missing and low or high values, using a semi-empirical formula instead.
@@ 1 Introduction
The octad (3500–5000 cm-1) and tetradecad regions (5000–6250 cm-1) in particular are crucial for satellite missions [lorente_methane_2021, malina_consistency_2022, chesnokova_estimation_2020], not only for its principal isotopologue but also for 13CH4 and CH3D, which are used to determine 13C/12C and D/H ratios. The octad has been improved in previous editions of HITRAN in the window observed by TROPOMI (4190–4340 cm-1). However, at higher wavenumbers of the octad, spectral parameters were of lower accuracy, incomplete, or both. Current and planned CH4 satellite sensors strive for  0.5% precision in the column (Ref. [jacob_quantifying_2022]; Table 1) for area flux mapping purposes, which requires similar accuracies for the intensities and widths and is not achieved in many regions. Even in the TROPOMI window, the HITRAN uncertainty code only reaches 4 (i.e., $0.0001\leq\Delta\nu<0.001$ cm-1 for line positions and $10\%\leq\frac{\Delta S_{ij}}{S_{ij}}<20\%$ for intensities).
Above, and including the tetradecad region, other problems arise from the lack of assignment and/or a value for the lower state energy ($E^{\prime\prime}$). Given the temperature range, these issues are limited in the Earth’s atmosphere, but make accurate spectral modeling impossible at more extreme temperatures (for example, low temperatures on Titan or high temperatures on hot-Jupiters and brown dwarfs). With that one should note, that while for higher temperatures one can use the HITEMP version of the database [10.3847/1538-4365/ab7a1a], it also has limitations due to difficulty calculating accurate line positions at higher wavenumbers. Furthermore, the origin of some of the parameters in HITRAN2020 is not well documented, especially for pressure-induced data such as those being the results of private communications. This can cause additional problems, preventing comparisons, or extrapolations and adjustments if the old data was generated by a model.
@@
The focus of the HITRAN2024 update was on the parameters in the 160-characters HITRAN format, while a few speed-dependent and first-order line-mixing parameters were added as well. Significant progress has been made between 900 and 14000 cm-1 [kefala_empirical_2024, nikitin_improved_2024, rodina_improved_2021, nikitin_preliminary_2013, starikova_assignment_2024, malarich_dual_2021, nikitin_first_2024, campargue_high_2023] for the three most abundant isotopologues, 12CH4, 13CH4, and 12CH3D, allowing for an update to the transition wavenumber, intensity, and assignments from the HITRAN2020 database using various recorded spectra for validation. These line lists also cover missing portions of the infrared spectrum that were previously challenging to obtain, as well as some in-between lines that were missing. Similar efforts have been put into obtaining good quality pressure-induced line-shape parameters [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023, dudaryonok_semiempirical_2018], with the differences being that we considered older work, with the intention of replacing the parameters for more lines than those in Section 2 to improve traceability. To this end, a combination of good quality measurements, with an empirical Padé-approximants model based on these measurements and semi-empirical models to fill the gaps was used. These results provide the necessary foundation for the inclusion of first-order and full line-mixing with speed-dependent profiles.
2 Additions and updates of non-pressure-induced data
@@
The newest versions of MARVELized-ExoMol [yurchenko_exomol_2024], TheoReTS [10.3847/1538-4357/aa8909], HITEMP [10.3847/1538-4365/ab7a1a] and MeCaSDa [MeCaSDa_2013], which is partly accessible via VAMDC portal [rixon_VAMDC_2011, albert_VAMDC_2020], line lists as well as lists from Nikitin et al. [nikitin_improved_2024, nikitin_preliminary_2013, nikitin_first_2024], Rodina et al. [rodina_improved_2021], Sung et al. [sung_new_2024], Starikova et al. [starikova_assignment_2024], Malarich et al. [malarich_dual_2021], and Campargue et al. [campargue_high_2023] were available for the HITRAN2024 update and are summarized for each region covered. These works were then validated against several high-quality, high-resolution FTS [2015ApJ...813...12H, bertin_co2_2024, birk_esa_2017, birk_measurement_2025] and CRDS spectra [birk_esa_2017] obtained under well-controlled laboratory conditions, when available for the updated region. The best parameters depending on the spectral region were selected and added to the HITRAN2024 database.
@@
Much of the focus was brought to the pentad, octad, and tetradecad regions, but many other spectral regions have also been updated for HITRAN2024. Changes from these lists were applied almost exclusively for the transition wavenumbers, intensities, and energy levels, since pressure-induced parameters of most of the available line lists were missing or approximate. Updates on these parameters are detailed in Sec. 3. The newly added lines, which were formerly missing in HITRAN, were added as is, but their pressure broadenings and shifts were replaced as well.
2.1 Full spectrum lists at disposal
Other compiled lists of calculated or measured line parameters were available for this update and summarized here. These were either included in other work or directly taken in some regions of HITRAN2024 as described starting Sec. 2.2.
@@ 2.1.1 ExoMol
The ExoMol project [tennyson_2024_2024] has produced a number of methane line lists of increasing completeness and accuracy designed for use at high temperatures [yurchenko_exomol_2014, yurchenko_hybrid_2017, yurchenko_exomol_2024]. The intensities in this list are calculated variationally. The line positions in the most recent of these [yurchenko_exomol_2024], known as the MM line list, was based on a comprehensive MARVEL (measured active vibration-rotation energy levels) study of 12CH4 by Kefala et al [kefala_empirical_2024]. This MARVEL study extracted 82173 rovibrational and rotational transitions from 96 literature sources containing high-resolution studies of 12CH4 spectra; the study identified a further 42 sources containing unassigned or partially assigned methane spectra which could not be included in the MARVEL study. An analysis of these sources is currently in progress and has so far led to the assignment of almost 10000 further lines, which extend the coverage of the previous study in terms of coverage of both rotational and vibrational states. This work can enable updates to future releases of both HITRAN and HITEMP. A HITRAN-format 296 K version of the MM line list generated using PyExoCross [Zhang_pyExoCross_2024] was used in this study, predominantly in the pentad region.
@@ 2.2.1 IAO FTIR measurements
2.2.2 HITRAN update in the 900–1050 cm-1 region
The new line list from Nikitin et al. [nikitin_improved_2024] (12CH4, 13CH4, 12CH3D & 13CH3D) completes the lower energy part of the dyad region, adding many lines including those for the lower-abundance 13CH3D isotopologue. Their data were merged directly to the region, replacing the transition wavenumbers for any difference ($\Delta\tilde{\nu}=0$ cm-1) between HITRAN2020 and the new parameters, the lower energy levels for $\Delta E^{\prime\prime}=0.1$ cm-1, and the intensities for $\Delta S_{ij}/S_{ij}=0.1$. About half of the lines in the list mainly served to update the positions and intensities with a few energy levels and assignments, while the other half was entirely new.
2.3 2500–3900 cm-1
2.3.1 HITRAN update in the 2500–3900 cm-1 region
Figure 1: Residuals obtained at 296 K between experiment from $S2$ Bertin & Vander Auwera [bertin_co2_2024] and Voigt calculations from the line lists used for the HITRAN2024 update
The residuals between $S1$, $S2$ and the models calculated using the HITRAN Application Programming Interface (HAPI) with lines from ExoMol (12CH4), MeCaSDa (12CH4 & 13CH4), and HITEMP (12CH4 & 13CH4) were used to compare the quality of each parameter on a line-by-line basis (these residuals are presended for $S2$ in Fig. 1). HITEMP was used when all other data were insufficient, since many lines are not fully assigned. By artificially increasing the pressure of the models, we could isolate the missing or strongly shifted lines in every line list, resulting in downward-pointing peaks. Using a combination of derivatives and thresholding, we could automatically find which areas were missing in HITRAN and which line lists could cover them. Because the residuals can be caused by multiple missing lines, every line above a certain intensity in a small window around the region to update was taken from the new line list. We then used the same technique, without the pressure increase, to find and update the weakly shifted lines in HITRAN. In addition to this automated approach, manual selections were made between lists for lines that were incorrectly updated or missed because of the various thresholds.
Using the same HITRAN2020 residuals, some lines were found to have a worse quality of fit than their counterparts back from HITRAN2000. These transition wavenumbers and intensities were manually reverted to their 2000 values.
2.4 3900–4900 cm-1
2.4.1 Kitt Peak & JPL FTS measurements of 13CH4
Multiple sets of 13CH4 line lists, either empirical or model predictions, were available from various sources, e.g., HITRAN 2020 [gordon_hitran2020_2022, brown_measurements_2016], JPL GFIT package (“atm.161”) [toon_spectrometric_2021], and Kitt Peak FTS (291 K) and JPl-FTS (80 K). More importantly, comprehensive updates to the previous work by Brown et al. [brown_measurements_2016] became available through the effective Hamiltonian modeling extended to the tetradecad region for hot band transitions. All of the line lists have been evaluated and cross-compared against high-resolution laboratory spectra obtained at 291 K and 80 K by generating synthetic spectra for the the same experimental conditions (e.g. sample pressure, temperature, gas cell pathlength, and spectral resolution) as the corresponding laboratory spectra, but using different sets of line lists available. These multiple sets of synthetic spectra were compared with lab spectra at each temperature to determine their quantum identifications and lower state energy values by adopting model calculation as presented in Sung et al. [sung_new_2024]
@@ 2.4.1 Kitt Peak & JPL FTS measurements of 13CH4
Table 1 presents statistics on the improvements (both additions and updates inclusive) as reported by Sung et al. [sung_new_2024], showing an almost ten-fold increase in the total number of line list entries in the 4000–4700 cm-1 wavenumber interval for 13CH4, when compared to the HITRAN2020.
@@
Figure 2: Comparison of synthetic spectra generated with HITRAN2020 (in pink) and our updated composite line list (in red) with high-resolution Kitt Peak FTS spectrum at 291 K (Left, in black) and JPL-FTS spectrum at 80 K (Right, in blue). This plots show better performance and broader spectral coverage by the new composite line list. Note that the synthetic spectra were computed for the same conditions as the corresponding laboratory spectra. Any H2O (not presented) and 12CH4 features (in green) were identified during the line-by-line matching, as presented for 12CH4 features.
@@
Finally, Fig. 2 presents a graphical summary of the updates, showing synthetic spectra generated with the HITRAN2020 line list and the composite line list reported in this work compared to two high-resolution laboratory FTIR spectra, respectively at 291 K( Left panel) and 80 K (Right panel). The synthetic spectra at 291 K and 80 K were computed for the same conditions as the corresponding laboratory spectra. It is shown that the composite line list covers much broader section of the Octad region as well as providing a better representation of the observed spectra at both temperatures thanks to the lower-state energy updates.
2.4.2 HITRAN update in the 3900–4900 cm-1 region
Spectra $S1$ and from Birk et al. [birk_esa_2017] ($S3$), including CRDS spectra not shown here, were used for the octad region. For the same reason that $S1$ shows a lower SNR, $S3$ was used for its window of 3900–4500 cm-1 instead.
Figure 3: Residuals obtained at 296 K and 0.17143 hPa between the experiment $S3$ from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 4: Residuals obtained at 198 K and 0.16941 hPa between the experiment $S3$ from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 5: Residuals obtained between experiment $S3$ from Birk and Voigt calculations from the past and new HITRAN versions
Unlike in the pentad region, several new line lists presented a clear improvement over large portions of the spectrum, as illustrated in Figures 3 and 4. This result alleviated the need for line-by-line analysis, with a few exceptions that were then fixed. In these large portions, the new lists were merged with HITRAN2020, adding new lines and updating existing ones unless no significant changes were made by conditions similar to those detailed in Sec. 2.2.2 , depending on the precision of the parameters. The residuals were then analyzed, keeping HITRAN2020 parameters if needed, ensuring that only the best parameters remained. This method was used for the data from Rodina et al. [rodina_improved_2021], between 4100 and 4190 cm-1, improving both CH4, 13CH4, and CH3D. Their line list extended to 4300 cm-1, but that region was previously updated in HITRAN2020 for TROPOMI [hu_operational_2016]. A list provided by Sung et al. for 13CH4 has been added as well between 4000–4700 cm-1, including the region between 4190–4300 cm-1 which replaces some lines, or intensities, incorrectly assigned to the main isotopologues. The list from Nikitin et al. [nikitin_preliminary_2013] improving CH3D between 4000–4554 cm-1 has been added, adding the lines missing in HITRAN2016 [gordon_hitran2016_2017] when the preliminary version of their data was first used. Finally, the empirical list from Toon (atm.161 [toon_atmospheric_2022]) replaced the region between 4341–4900 cm-1 for lines above the intensity of $10^{-27}$ cm//molecule, improving CH4, 13CH4, and CH3D.
We observed an approximately systematic shift in the transition wavenumber of lines from Daumont et al. [daumont_new_2013]. Applying a constant shift of 0.001 cm-1 appeared to improve the residuals, particularly below 4200 cm-1. The effect is limited above 4375 cm-1 as the addition of atm.161 list already improved that region appreciably. The final residuals for the latest HITRAN versions are shown in Figure 5, confirming the great improvement above 4375 cm-1, after keeping the region unchanged in the last 3 versions.
2.5 4970–5300 cm-1
2.5.1 IAO analyses of JPL FTS measurements of 13CH4
@@
2.5.2 HITRAN update in the 4970–5300 cm-1 region
2.6 5900–6530 cm-1
2.6.1 NIST, DLR & USTC measurements
The tetradecad spectral region has been the subject of extensive study, including the WKLMC empirical line list [campargue_2013], the GOSAT 2014 line list [nikitin_gosat-2014_2015] (which was the basis of both HITRAN2016 and HITRAN2020 for this spectral region), the widely used empirical line list of Devi 2015 [devi_self-_2015], and the Nikitin 2017 analysis of WKLMC [nikitin_analysis_2017]. Significant differences exist between these line lists, complicated by correlation between parameters in these dense and complex spectra and the difficulty in identifying and assigning weaker transitions. These complications lead to line lists that trade parameter accuracy for self-consistency, thus impeding efforts to refine line parameters.